# skipabove=5pt, innertopmargin=0pt

### v1

@@ 7 Conclusion
The core problem we address in this paper is the role of reasoning in AI safety. While there have been recent efforts arguing that replacing reasoning with increased compute can lead to better defense mechanisms, these approaches contain a fundamental oversight. They fail to consider that attackers may also leverage reasoning to bypass guardrails. This paper defines adversarial reasoning, demonstrates a practical implementation, and provides state-of-the-art results on attack success rate.
@@
Our work points to new directions for understanding and improving language model security. By bridging reasoning frameworks with adversarial attacks, we have demonstrated how structured exploration of the prompt space can reveal vulnerabilities even in heavily defended models. This suggests that future work on model alignment may need to consider not just individual prompts but entire reasoning paths when developing robust defenses. The success of our transfer attack methodology also highlights the importance of considering multiple surrogate models when evaluating model security. Looking ahead, our findings point to several promising research directions, including developing more sophisticated reasoning-guided search strategies, exploring hybrid approaches that combine token-level and semantic-level optimization, and investigating how process-based rewards could be incorporated into defensive training. Finally, while our study has focused on textual LLMs, our framework can potentially be relevant to the broader class of LLM-driven agents [ASD+24]. In particular, our methods can be naturally extended to LLM-controlled robots [LHX+23, KNC+23, VBBK24], web-based agents [WSK+24], and AI-powered search engines [RBC+24]. Recent work [RRK+24] underscores this connection by demonstrating that vulnerabilities identified in textual models can be transferred to real-world scenarios.
Appendix A Experiments setting
For running Algorithm 1 in Section 5.1, we loaded the target models on our local GPUs to read to the log-probs, but used TogetherAI [Tog] for collecting the full responses. We also used TogetherAI for the attacker, Feedback LLM, and Refiner LLM in all the sections. We used one NVIDIA A100 for our experiments.
@@ Appendix A Experiments setting
We chose to put GPT4o and Llama-3.1-405B in the black-box category. Despite the previous attempt to extract the entire log-prob vector based on the top-5 log-probs for GPT-4 [HBC+24], there is no guarantee that OpenAI will preserve this feature in later releases, so we have included this model as one of the black-box models. For Llama-3.1-405B, we used TogetherAI for querying since the model would not fit to our GPUs, and TogetherAI does not give access to the entire log-prob vector.
@@
Appendix B Additional experiments
@@
For some of the most common LLMs, we illustrate their vulnerabilities in different categories of Harmbench [MPY+24]. Figure 6 shows that Claude has a stronger performance (lower ASR) compared to other models except in the "cyber_intrusion" category where the model is outperformed by OpenAI o1-preview and Gemini-1.5-pro.
@@
Appendix C System Prompts
@@
We get the judge provided by Harmbench [MPY+24] to give a verdict for the jailbreaking attempts. This model is fine-tuned based on Llama-2-13B [AI@23] with a knowledge distillation process from GPT-4, and a meticulously crafted prompt for querying the models. Below we have attached the prompt used for this manner, which is a slightly modified version of Harmbench’s original prompt. This prompt emphasizes on detailed-oriented outputs, precisely addressing what the original intent demands. Consequently, if the output is malign but irrelevant to the demand, falls short of the request (e.g., has not generated the code, or lacks the steps where it has to be step-by step), or they it is vague and not a useful answer, the judge will it them by default.
@@
We use a system prompt that instructs the attacker model with a jailbreaking state without providing any in-context examples. While in-context learning has been shown to improve performance in many scenarios [DLD+24], it also biases the model toward specific attack strategies demonstrated in the examples. In contrast, an ideal search should autonomously identify the most effective attack method for a given target. By omitting in-context examples, our approach reduces reliance on the system prompt—ensuring that changes in the prompt do not affect effectiveness, as no additional instructions are embedded. To structure the attacker’s output, we explicitly require two components: ‘Thoughts’ and ‘Prompt’. Consistent with prior work [CRD+24, MZK+24], we find that prompting the model to articulate its reasoning in ‘Thoughts’ before generating an attack prompt improves adherence to instructions and responsiveness to feedback. Finally, we use the generated text in ‘Prompt’ directly for jailbreaking.
@@
Our initial experiments showed that when loss values were attached to their corresponding attacking prompts and presented in a random order, weaker LLMs such as Vicuna-13B-v1.5 and Mixtral-8x7B-v0.1 struggled to correctly sort them based on loss values. We believe this confusion arises due to long text segments appearing between the loss values, leading to a length generalization issue [LLH+23, AWA+22]. This suggests that LLMs cannot robustly utilize information when the context is too long and tend to neglect text in the middle [LLH+23]. To mitigate this limitation, we pre-sort the prompts according to their loss values and present them to the LLM without explicitly attaching the corresponding losses. Second, we explicitly instruct the model to identify patterns in the attacking prompts that contribute to increasing loss values. We require the model to compare arbitrary pairs of prompts to analyze differences in their attacking strategies. This approach encourages the model to leverage all available attacking prompts for feedback generation.
@@
As pointed out in [YBB+24], the functionality of the Refiner LLM is akin to applying the gradients in SGD. The primary role of the Refiner LLM is to read the feedback and apply its key points to the reasoning string, and at the same time, neglect those part that are not helpful (e.g., some parts of the feedback might be already included in the reasoning string). Hence, the system prompt for the Refiner LLM is designed to mask out these unwanted artifacts. Another important aspect is to apply the these steps incrementally; The output of this function must be close to the input, and semantically resemble the original reasoning string with few modification. This will make the search method stable in the prompt space, where one step does not result in abrupt changes in the attacking prompt.
@@ Appendix D Examples
Figure 10: A full example of exploring the prompt-space for 6 iterations and jailbreaking OpenAI o1. Note that each node has $n/k=8$ children, where only the explored ones are shown in this graph for simplicity.
w
Figure 11: An example of exploring the prompt-space for only 1 iteration that jailbreaks Claude. Note that the initial node has 8 children where only the explored one is shown here for simplicity.

### v2

@@ 6 Conclusion
This paper investigates the role of reasoning in AI safety, showing that defenses that simply trade reasoning for more compute overlook the fact that attackers may also leverage reasoning to bypass guardrails. This paper defines adversarial reasoning, demonstrates a practical implementation, and provides state-of-the-art results on attack success rate.
@@
Our work points to new directions for understanding and improving language model security. By bridging reasoning frameworks with adversarial attacks, we have demonstrated how structured exploration of the prompt space can reveal vulnerabilities even in heavily defended models. This suggests that future work on model alignment may need to consider not just individual prompts but entire reasoning paths when developing robust defenses. The success of our transfer attack methodology also highlights the importance of considering multiple surrogate models when evaluating model security. Looking ahead, our findings point to several promising research directions, including developing more sophisticated reasoning-guided search strategies, exploring hybrid approaches that combine token-level and semantic-level optimization, and investigating how process-based rewards could be incorporated into defensive training. Finally, while our study has focused on textual LLMs, our framework can potentially be relevant to the broader class of LLM-driven agents (Andriushchenko et al., 2024b). In particular, our methods can be naturally extended to LLM-controlled robots (Liang et al., 2023; Karamcheti et al., 2023; Vemprala et al., 2024), web-based agents (Wu et al., 2024), and AI-powered search engines (Reuel et al., 2024). Recent work (Robey et al., 2024a) underscores this connection by demonstrating that vulnerabilities identified in textual models can be transferred to real-world scenarios.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Additional Related Work
Token-space Jailbreaking. Token-space attacks (Shin et al., 2020; Wen et al., 2023; Zou et al., 2023; Hayase et al., 2024; Andriushchenko et al., 2024a) modify the input at the token level to decrease some loss value. For example, the GCG algorithm (Zou et al., 2023), one of the first transferrable token-level attacks to achieve significant success rates on aligned models, uses the gradient of the loss to guide the greedy search. Subsequent work has refined this approach to obtain lower computational cost and improved effectiveness (Liao & Sun, 2024; Jia et al., 2024), including token-level modifications by other heuristics and without a gradient (Hayase et al., 2024) and random searches over cleverly chosen initial prompts (Andriushchenko et al., 2024a). We adopt the use of a loss function from these methods as a signal to inform how to navigate the prompt space for better jailbreaks while remaining gradient-free.
Semantic-space Jailbreaking. These methods often rely on a "red-teaming" LLM to generate adversarial prompts (Perez et al., 2022; Wei et al., 2023a; Sadasivan et al., 2024; Chao et al., 2024b; Liu et al., 2024b; Mehrotra et al., 2024; Zeng et al., 2024; Samvelyan et al., 2024; Liu et al., 2024a). Methods such as PAIR (Chao et al., 2024b) deploy a separate LLM, called the "attacker", which uses a crafted system prompt to interact with the target LLM over multiple rounds and generate semantic jailbreaks; they operate in a black-box manner, requiring only the target’s outputs, and are highly transferrable (Chao et al., 2024b). Some other methods fine-tune a model to generate the attacking prompts (Perez et al., 2022; Ge et al., 2023; Zeng et al., 2024; Paulus et al., 2024; Beetham et al., 2024), though this demands substantial computational resources. Rather than fine-tuning, we rely on increased test-time computation (Snell et al., 2024), while others start from expert-crafted prompts (e.g., DAN) and refine them via genetic algorithms (Liu et al., 2024b; Samvelyan et al., 2024; Lapid et al., 2024). Like these methods, our approach generates semantically meaningful jailbreaks by using another LLM as the attacker, however, our approach is significantly different from the prior work as we develop reasoning modules based on the loss values to better navigate the prompt space.
Appendix B Experiments setting
Human evaluation
We use the HarmBench judge (Mazeika et al., 2024) to evaluate the target responses. However, as explained in Section 5, we manually verify all of the jailbreaks marked as positive by the judge. We remark that the additional human-based evaluation on the top of the HarmBench judge is in fact necessary, and has been done in previous work (e.g., (Zou et al., 2024). This is because the HarmBench judge occasionally makes mistakes by detecting harmless answers as jailbreaks. In an attempt to keep the manual evaluation impartial, we enlisted three experts in jailbreaking to evaluate the responses according to the same outline given to the Harmbench judge (the system prompt provided in the appendix). Experts do not know the algorithms and are only provided with the tasks, jailbreaking prompts, and responses. We classify a response as a jailbreak only if all three experts unanimously agree.
@@ Appendix B Experiments setting
We chose to put GPT4o and Llama-3.1-405B in the black-box category. Despite the previous attempt to extract the entire log-prob vector based on the top-5 log-probs for GPT-4 (Hayase et al., 2024), there is no guarantee that OpenAI will preserve this feature in later releases, so we have included this model as one of the black-box models. For Llama-3.1-405B, we used TogetherAI for querying since the model would not fit to our GPUs, and TogetherAI does not give access to the entire log-prob vector.
@@
Compute
For running Algorithm 1 in Section 5.1, we loaded the target models on our local GPUs to read to the log-probs, but used TogetherAI for collecting the full responses. We also used TogetherAI for the attacker, Feedback LLM, and Refiner LLM in all the sections. We used one NVIDIA A100 for our experiments.
@@
Appendix C Additional experiments
@@
For some of the most common LLMs, we illustrate their vulnerabilities in different categories of Harmbench (Mazeika et al., 2024). Figure 6 shows that Claude has a stronger performance (lower ASR) compared to other models except in the "cyber_intrusion" category where the model is outperformed by OpenAI o1-preview and Gemini-1.5-pro.
@@
Appendix D System Prompts
@@
We get the judge provided by Harmbench (Mazeika et al., 2024) to give a verdict for the jailbreaking attempts. This model is fine-tuned based on Llama-2-13B (Meta AI, 2023) with a knowledge distillation process from GPT-4, and a meticulously crafted prompt for querying the models. Below we have attached the prompt used for this manner, which is a slightly modified version of Harmbench’s original prompt. This prompt emphasizes on detailed-oriented outputs, precisely addressing what the original intent demands. Consequently, if the output is malign but irrelevant to the demand, falls short of the request (e.g., has not generated the code, or lacks the steps where it has to be step-by step), or they it is vague and not a useful answer, the judge will it them by default.
@@
We use a system prompt that instructs the attacker model with a jailbreaking state without providing any in-context examples. While in-context learning has been shown to improve performance in many scenarios (Dong et al., 2024), it also biases the model toward specific attack strategies demonstrated in the examples. In contrast, an ideal search should autonomously identify the most effective attack method for a given target. By omitting in-context examples, our approach reduces reliance on the system prompt—ensuring that changes in the prompt do not affect effectiveness, as no additional instructions are embedded. To structure the attacker’s output, we explicitly require two components: ‘Thoughts’ and ‘Prompt’. Consistent with prior work (Chao et al., 2024b; Mehrotra et al., 2024), we find that prompting the model to articulate its reasoning in ‘Thoughts’ before generating an attack prompt improves adherence to instructions and responsiveness to feedback. Finally, we use the generated text in ‘Prompt’ directly for jailbreaking.
@@
Our initial experiments showed that when loss values were attached to their corresponding attacking prompts and presented in a random order, weaker LLMs such as Vicuna-13B-v1.5 and Mixtral-8x7B-v0.1 struggled to correctly sort them based on loss values. We believe this confusion arises due to long text segments appearing between the loss values, leading to a length generalization issue (Liu et al., 2023; Anil et al., 2022). This suggests that LLMs cannot robustly utilize information when the context is too long and tend to neglect text in the middle (Liu et al., 2023). To mitigate this limitation, we pre-sort the prompts according to their loss values and present them to the LLM without explicitly attaching the corresponding losses. Second, we explicitly instruct the model to identify patterns in the attacking prompts that contribute to increasing loss values. We require the model to compare arbitrary pairs of prompts to analyze differences in their attacking strategies. This approach encourages the model to leverage all available attacking prompts for feedback generation.
@@
As pointed out in (Yuksekgonul et al., 2024), the functionality of the Refiner LLM is akin to applying the gradients in SGD. The primary role of the Refiner LLM is to read the feedback and apply its key points to the reasoning string, and at the same time, neglect those part that are not helpful (e.g., some parts of the feedback might be already included in the reasoning string). Hence, the system prompt for the Refiner LLM is designed to mask out these unwanted artifacts. Another important aspect is to apply the these steps incrementally; The output of this function must be close to the input, and semantically resemble the original reasoning string with few modification. This will make the search method stable in the prompt space, where one step does not result in abrupt changes in the attacking prompt.
@@ Appendix E Examples
Figure 10: A full example of exploring the prompt space for 6 iterations and jailbreaking OpenAI o1. Note that each node has $n/k=8$ children, where only the explored ones are shown in this graph for simplicity.
Figure 11: An example of exploring the prompt space for only 1 iteration that jailbreaks Claude. Note that the initial node has 8 children where only the explored one is shown here for simplicity.